OpenStudy (anonymous):

Shelly sews a blanket that has an area of 170 square feet. It has 30 square blocks, each the same size. What is the approximate length of each side of a block?

OpenStudy (nubeer):

let each side of small square =x so area of small square = x*x =x^2.. agreed>?

OpenStudy (fellowroot):